# First Digit Distributions of Gamma-Ray Bursts

### v1

@@ Abstract
The occurrence of the first significant digits from real world sources is usually not equally distributed, but is consistent with a logarithmic distribution instead, known as Benford’s law. In this work, we perform a comprehensive investigation on the first digit distributions of the duration, fluence, and energy flux of gamma-ray bursts (GRBs) for the first time. For a complete GRB sample, we find that the first digits of the duration and fluence adhere to Benford’s law. However, the energy flux shows a significant departure from this law, which may be due to the fact that a considerable part of the energy flux measurements are restricted by lack of spectral information. Based on the conventional duration classification scheme, we also check if the durations and fluences of long and short GRBs (with duration s and s, respectively) obey Benford’s law. We find that the fluences of both long and short GRBs still agree with the Benford distribution, but their durations do not follow Benford’s law. Our results hint that the long–short GRB classification scheme does not directly represent the intrinsic physical classification scheme.
@@ 3 Summary
In this work, we have performed a systematic investigation on the first digit distributions of the duration, fluence, and energy flux of GRBs. For the complete GRB sample, our results show that the first digits of the duration and fluence are not uniformly distributed, but that small digits are more common than big ones according to a logarithmic distribution as expected by Benford’s law. However, not all quantities obey Benford’s law. Artificial and restricted quantities often deviate from the law, such as the energy fluxes of GRBs in our study. The main reason for the deviation may be that the energy flux measurements are restricted by spectral fits for each GRB, and only 63% of bursts have the required spectral information. Thus, there is not enough dynamic range to make the energy fluxes be fully compliant with Benford’s law.
The Fermi-GBM burst catalog also revealed a bimodal distribution, and two classes of GRBs, i.e., long versus short GRBs with a separation at about 2 s, were confirmed (Kouveliotou et al. 1993; von Kienlin et al. 2020). Using the conventional long–short GRB classification scheme, we found that there were long GRBs (with s) and short GRBs (with s). Here, the first digits of the duration and fluence of long and short GRBs were also examined for adherence to Benford’s distribution. We found that the fluences of both long and short GRBs still follow Benford’s law, but their durations are no longer consistent with this law. Our results indicate that the data of fluence seems to be very natural and believable, but that is not always a good quantity to conduct GRB classification. That is, the long–short GRB classification scheme does not directly represent the intrinsic physical classification scheme.
@@ Acknowledgements.
This work is partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (grant No. XDB0550400), the National Natural Science Foundation of China (grant Nos. 12373053 and 12321003), the Key Research Program of Frontier Sciences (grant No. ZDBS-LY-7014) of Chinese Academy of Sciences, and the Natural Science Foundation of Jiangsu Province (grant No. BK20221562).References

### v2

@@ Abstract
The occurrence of the first significant digits from real world sources is usually not equally distributed, but is consistent with a logarithmic distribution instead, known as Benford’s law. In this work, we perform a comprehensive investigation on the first digit distributions of the duration, fluence, and energy flux of gamma-ray bursts (GRBs) for the first time. For a complete GRB sample detected by the Fermi satellite, we find that the first digits of the duration and fluence adhere to Benford’s law. However, the energy flux shows a significant departure from this law, which may be due to the fact that a considerable part of the energy flux measurements are restricted by lack of spectral information. Based on the conventional duration classification scheme, we also check if the durations and fluences of long and short GRBs (with duration s and s, respectively) obey Benford’s law. We find that the fluences of both long and short GRBs still agree with the Benford distribution, but their durations do not follow Benford’s law. Our results hint that the long–short GRB classification scheme does not directly represent the intrinsic physical classification scheme.
@@ 3 Summary
In this work, we have performed a systematic investigation on the first digit distributions of the duration, fluence, and energy flux of GRBs. For the complete GRB sample detected by Fermi-GBM, our results show that the first digits of the duration and fluence are not uniformly distributed, but that small digits are more common than big ones according to a logarithmic distribution as expected by Benford’s law. However, not all quantities obey Benford’s law. Artificial and restricted quantities often deviate from the law, such as the energy fluxes of GRBs in our study. The main reason for the deviation may be that the energy flux measurements are restricted by spectral fits for each GRB, and only 63% of bursts have the required spectral information. Thus, there is not enough dynamic range to make the energy fluxes be fully compliant with Benford’s law.
The Fermi-GBM burst catalog also revealed a bimodal distribution, and two classes of GRBs, i.e., long versus short GRBs with a separation at about 2 s, were confirmed (Kouveliotou et al. 1993; von Kienlin et al. 2020). Using the conventional long–short GRB classification scheme, we found that there were long GRBs (with s) and short GRBs (with s). Here, the first digits of the duration and fluence of long and short GRBs were also examined for adherence to Benford’s distribution. We found that the fluences of both long and short GRBs still follow Benford’s law, but their durations are no longer consistent with this law. Our results indicate that the data of fluence seems to be very natural and believable, but that is not always a good quantity to conduct GRB classification. That is, the long–short GRB classification scheme does not directly represent the intrinsic physical classification scheme (Zhang et al. 2007; Lü et al. 2010; Qin et al. 2013).
To cross-check of the results produced on Fermi-GBM, we also analyzed the first digit distributions for the GRB data observed with the Burst Alert Telescope (BAT) onboard the Swift satellite (please see Appendix A for more details). We found that (i) the derived -values change quantitatively, though the qualitative results and conclusions remain the same for the duration and fluence distributions of the overall GRB data, independent of what kind of GRB mission is considered; (ii) Benford’s law is still followed by the fluence distributions of both long and short GRBs, for each GRB mission; (iii) the duration distributions of the long and short GRB groups observed with Swift-BAT are generally consistent with Benford’s law, but those of the Fermi-GBM sample deviate from this law. Since the duration distributions of long and short GRBs for samples observed with different missions do not always apply to Benford’s law, we emphasize agian that the duration classification scheme does not always match the intrinsic physical classification scheme.
@@ Acknowledgements.
We are grateful to the anonymous referee for helpful comments. This work is partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (grant No. XDB0550400), the National Natural Science Foundation of China (grant Nos. 12373053 and 12321003), the Key Research Program of Frontier Sciences (grant No. ZDBS-LY-7014) of Chinese Academy of Sciences, and the Natural Science Foundation of Jiangsu Province (grant No. BK20221562).References
@@
- Zhang et al. (2007) Zhang, B., Zhang, B.-B., Liang, E.-W., et al. 2007, ApJ, 655, L25
Appendix A First digit distributions for Swift-BAT GRBs
To cross-check the results produced on Fermi-GBM, we also analyze the first digit distributions for the GRB data observed with Swift-BAT. For the Swift-BAT sample, the durations (in units of s) and fluences (in units of erg ) in the 15–150 keV energy range are taken from the online burst catalog.222https://swift.gsfc.nasa.gov/archive/grb_table/ The dataset contains 1627 GRBs up to 2024 January. We remove 143 bursts for which no duration or fluence measurements were available. In total, there are 1484 GRBs (including 1354 long bursts with s and 130 short ones with s) for us to perform the first digit analyses.
The first digit distributions of the duration and fluence for all 1484 GRBs, 1354 long GRBs, and 130 short GRBs are illustrated in Figures 5, 6, and 7, respectively. A tabular summary of our Benford analyses for the Swift-BAT sample can be found in Table 2. The Pearson tests and the corresponding -values indicate that the durations and fluences for the overall sample and the subsamples of long and short GRBs are all roughly consistent with Benford’s law. As shown in Figure 7, the digits 2, 5, and 9 are smaller than the expected Benford distributions. That is, the duration and fluence of short GRBs do not seem to fit Benford’s law through the eyes, but their corresponding -values support they do. The optical illusion may be caused by the relatively small sample size of short GRBs.
The comparison between the Fermi-GBM and Swift-BAT samples may be summarized as follows: (i) the first digit distributions of the duration and fluence of the overall GRB data conform to Benford’s law, independent of what kind of GRB mission is considered; (ii) Benford’s law is still followed by the fluence distributions of both long and short GRBs, for each GRB mission; (iii) the duration distributions of the long and short GRB groups observed with Fermi-GBM do not obey Benford’s law, but those of the Swift-BAT sample are generally consistent with this law. Note that the ratios of short-to-long GRB numbers with a division of s for the Fermi-GBM and Swift-BAT samples are 603:3061 (1:5.1) and 130:1354 (1:10.4), respectively. Obviously, the distribution is instrument dependent. Again, our results suggest that the duration classification scheme does not always match the intrinsic physical classification scheme. It it does, then the duration distributions of both long and short GRBs for samples observed with different missions should always adhere to Benford’s law.
Dataset | Physical quantity | Number | -value | |
All GRBs | Duration | 1484 | 16.4/8 | 0.04 |
Fluence | 1484 | 8.3/8 | 0.41 | |
Long GRBs | Duration | 1354 | 12.7/8 | 0.12 |
Fluence | 1354 | 7.5/8 | 0.48 | |
Short GRBs | Duration | 130 | 10.0/8 | 0.26 |
Fluence | 130 | 8.3/8 | 0.41 |